What is the variance of the returns on a security given the following information?

State of the Economy Probability E(R)
Boom 0.25 16 %
Normal 0.45 10 %
Recession 0.30 -8 %

Multiple Choice

  • 91.35

  • 72.87

  • 64.38

  • 48.18

  • 56.23
Answer #1

Answer: Option A:91.35

Expected Return =

Expected return E(R)= 0.25*16%+0.45*10%+0.3*-8%=6.10%

State of economy Probability P E(A)
Boom 0.25 16%
Normal 0.45 10%
Recession 0.3 -8%
Expected return E(R) 6.10%

So Variance V=0.25*(16%-6.10%)^2+0.45*(10%-6.10%)^2+0.3*(-8%-6.10%)^2=91.35